# SuperGrasp: Single-View Object Grasping via Superquadric Similarity Matching, Evaluation, and Refinement

### 31 Mar 2026 v1

@@ V Experiments
We conduct extensive comparative experiments in both simulation and real-world environments to demonstrate that: 1) the proposed superquadric similarity matching module can stably generate rich and effective grasp candidates; 2) the proposed evaluation and refinement module can accurately evaluate and refine grasp poses; and 3) the overall framework exhibits strong generalization capability and real-time performance. We choose PointNetGPD[12] and Contact-GraspNet[24] as baselines. For a fair comparison among different methods, we adopt the following evaluation metrics:
@@
Ours (S1) + PointNetGPD[12] 63.34% 74.30% 341 38.42% 47.83% 747 31.62% 39.81% 351
PointNetGPD[12] + Ours (S2) 93.60% 55.26% 172 79.02% 48.33% 367 87.85% 33.21% 107
Ours w/o RNet 95.29% 87.67% 276 87.56% 84.50% 579 85.57% 85.00% 298
Ours 95.08% 95.00% 285 92.82% 92.67% 599 89.03% 94.67% 319
@@ V-A Simulation Experiments
We construct a $50\,\mathrm{cm}\times 50\,\mathrm{cm}$ grasping workspace in the PyBullet[21] simulation environment and deploy a fixed depth camera to capture scene point clouds. To thoroughly evaluate the performance of the proposed method, we design three experimental settings: 1) scenes containing 10 objects (moderately cluttered), 2) scenes containing 20 objects (densely cluttered), and 3) scenes containing 10 unseen objects, as illustrated in Fig. 6. For a fair comparison, the same random seeds are used for all methods in each setting to ensure identical scene configurations. Each setting is tested over 30 trials. The averaged results are reported in Table I.
As shown in Table I, in the relatively cluttered scene with 10 objects, our method achieves both an average grasp success rate and an average task success rate of over $95\%$, demonstrating its strong stability and efficiency. Compared with the structurally similar two-stage method PointNetGPD, our method improves the average grasp success rate by $16.32\%$ and the average task success rate by $64.59\%$. Compared with the end-to-end Contact-GraspNet baseline, our method improves the average grasp success rate by $16.83\%$ and the average task success rate by $14.67\%$. These results jointly validate the superiority of the proposed two-stage framework. In the more challenging setting with 20 objects, our method still maintains both the average grasp success rate and the average task success rate above $92\%$. Specifically, it surpasses PointNetGPD by $33.39\%$ and Contact-GraspNet by $17.40\%$ in grasp success rate, while improving the task success rate by $58.00\%$ and $9.84\%$, respectively. These results demonstrate the robustness and generalization capability of our method in highly cluttered environments. When evaluated on unseen objects that are not included in the training set, our method still achieves an average task success rate above $94\%$. It outperforms PointNetGPD by $87.63\%$ and Contact-GraspNet by $25\%$ in task success rate, while the average grasp success rate is improved by $55.13\%$ and $25.12\%$, respectively. These results further verify the strong object-level generalization ability of the proposed framework.
@@ V-B Ablation Study
We conduct the ablation experiments in the same PyBullet simulation environment under three settings: 1) scenes with 10 objects, 2) scenes with 20 objects, and 3) scenes with 10 unseen objects. Each setting is tested over 30 trials, and the averaged results are reported in Table I. When our Stage-1 grasp candidate generation module is combined with the downstream grasp evaluation module of PointNetGPD, the task success rate is significantly improved compared with PointNetGPD in all settings. Specifically, the average task success rate is improved by $43.89\%$ in the 10-object setting, $13.16\%$ in the 20-object setting, and $32.77\%$ in the unseen-object setting. These results confirm that our sampling strategy can stably provide rich and effective grasp candidates, thereby improving the overall task success rate. However, the average grasp success rate decreases in all settings. We attribute this to the increased number of grasp attempts together with the limited accuracy of the PointNetGPD evaluation module. Notably, when the grasp candidate generation module of PointNetGPD is combined with our Stage-2 module, i.e., grasp evaluation and refinement, the average grasp success rate is substantially improved in all three settings. In particular, in the unseen-object setting, the grasp success rate is improved by $53.95\%$. Overall, the average task success rate is improved by $21.56\%$. These results strongly validate the accuracy, robustness, and generalization capability of our evaluation and refinement module. In addition, we observe that the refinement branch has only a limited effect on grasp success rate in the 10-object setting, but brings clear improvements in the more cluttered 20-object setting and in the unseen-object setting with more complex surface geometry. Moreover, the average task success rate across the three settings is further improved by $8.39\%$, which verifies the necessity and effectiveness of the proposed refinement strategy.
@@ V-C Real-World Experiments
To further validate the strong generalizability of the proposed method, we construct a real-world grasping test set consisting of 30 similar objects and 25 novel objects whose shapes and sizes are substantially different from those in the training set. We design two cluttered real-world scenarios for evaluation: 1) randomly placing 10 similar objects and 2) randomly placing 10 novel objects. Each scenario is tested over 10 trials, and some representative scenarios are shown in Fig. 7. The hardware setup is as follows. We use a xMate ER7Pro robotic arm, an AG-105-145 gripper, and a ZED 2i camera. The entire system is integrated through ROS, and inverse kinematics is solved using MoveIt![4]. We employ FoundationStereo [28] to estimate depth from stereo RGB images and reconstruct the scene point cloud. We then use SAM [9] to segment the target object and obtain its point cloud. The experimental results are reported in Table II.
@@
Ours 0.356 0.433 0.789
As shown in Table II, in the similar-object setting, our method achieves both a grasp success rate and a task success rate of $98\%$. Specifically, the average grasp success rate exceeds that of PointNetGPD by $39.18\%$ and Contact-GraspNet by $19.05\%$, while the average task success rate is higher than that of PointNetGPD by $78\%$ and that of Contact-GraspNet by $38\%$. In the unseen-object setting, our method still maintains strong performance, with both the grasp success rate and the task success rate remaining above $94\%$. In this setting, the average grasp success rate exceeds PointNetGPD by $41.74\%$ and Contact-GraspNet by $19.47\%$, while the average task success rate is higher than that of PointNetGPD by $85\%$ and that of Contact-GraspNet by $43\%$. These results demonstrate the stability and strong generalization capability of the proposed method. During the experiments, we also record the average time cost of each stage, as reported in Table III. The time consumption of each stage in our method is lower than that of PointNetGPD, especially in the grasp candidate generation stage, although the overall runtime is slightly higher than that of Contact-GraspNet.
In addition, we observe that the primary cause of grasp failure in real-world experiments is self-occlusion of the target object or occlusion caused by surrounding objects, which introduces unknown space in the scene and eventually causes collisions between the gripper and the object during execution.
@@ VI Conclusion
We propose SuperGrasp, a novel two-stage grasping framework which addresses the challenge of grasp sampling under incomplete geometric information through superquadric representation and similarity matching in the first stage. In the second stage, we expand the grasp perception range and utilize the initial gripper closure region as an anchor region to evaluate and refine grasp candidates. Extensive experiments demonstrate the effectiveness of each stage, and the overall framework achieves strong performance with robust generalization across diverse scenarios. Although the real-time performance of each stage of the grasping process is better than that of the two-stage PointNetGPD method, the cumulative time consumption of each stage is slightly higher than that of the end-to-end method Contact-GraspNet. In the future, we will continue to study how to improve the real-time performance of grasping while maintaining strong generalization and stability.

### 15 Apr 2026 v2

@@ V Experiments
We conduct extensive comparative experiments in both simulation and real-world environments to demonstrate that: 1) the proposed superquadric similarity matching module can stably generate rich and effective grasp candidates; 2) the proposed evaluation and refinement module can accurately evaluate and refine grasp poses; and 3) the overall framework exhibits good generalization capability and real-time performance. We selected PointNetGPD[13], Contact-GraspNet [24], and Chen et al.[3] as baselines, where Chen et al.’s method was evaluated only on novel objects since it is designed specifically for novel object settings. Evaluation metrics are as follows:
@@
Chen et al.[3] 58.14% 83.33% 430
Ours (S1) + PointNetGPD[13] 63.34% 72.00% 341 38.42% 47.83% 747 31.62% 37.00% 351
PointNetGPD[13] + Ours (S2) 93.60% 53.67% 172 79.02% 48.33% 367 87.85% 31.33% 107
Ours w/o RNet 95.29% 87.67% 276 87.56% 84.50% 579 89.56% 88.67% 297
Ours 94.02% 94.33% 301 92.82% 92.67% 599 89.03% 94.67% 319
@@ V-A Simulation Experiments
We use the PyBullet[5] simulation environment and deploy a fixed depth camera to capture point clouds in the scene. To thoroughly evaluate the performance of the proposed method, we design three experimental settings: 1) scenes containing 10 objects (moderately cluttered), 2) scenes containing 20 objects (densely cluttered), and 3) scenes containing 10 unseen objects (i.e., objects not included in the training dataset), as illustrated in Fig. 6. For a fair comparison, the same random seeds are used for all methods in each setting to ensure identical scene configurations. Each setting is tested over 30 trials. The average results are reported in Table I.
As shown in Table I, in the relatively cluttered scene with 10 objects, our method achieves both an average grasp success rate and an average task success rate of over $94\%$, demonstrating its strong stability and efficiency. Compared with the structurally similar two-stage method PointNetGPD, our method improves the average grasp success rate by $15.26\%$ and the average task success rate by $64.66\%$. Compared with the end-to-end Contact-GraspNet baseline, our method improves the average grasp success rate by $15.77\%$ and the average task success rate by $14\%$. These results jointly validate the superiority of the proposed two-stage framework. In the more challenging setting with 20 objects, our method still maintains both the average grasp success rate and the average task success rate above $92\%$. Specifically, it surpasses PointNetGPD by $33.39\%$ and Contact-GraspNet by $17.4\%$ in grasp success rate, while improving the task success rate by $58\%$ and $9.84\%$, respectively. These results demonstrate the robustness and generalization capability of our method in highly cluttered environments. When evaluated on unseen objects that are not included in the training set, our method still achieves an average task success rate above $94\%$. It outperforms PointNetGPD by $88\%$, Contact-GraspNet by $25\%$, and Chen et al. by $11.34\%$ in task success rate, while the average grasp success rate is improved by $55.13\%$, $25.12\%$, and $30.89\%$, respectively. These results further verify the proposed framework’s good object-level generalization.
@@ V-B Ablation Study
We conduct the ablation experiments in the same PyBullet simulation environment under three settings: 1) scenes with 10 objects, 2) scenes with 20 objects, and 3) scenes with 10 unseen objects. Each setting is tested over 30 trials, and the averaged results are reported in Table I. When our Stage-1 grasp candidate generation module is combined with the PointNetGPD downstream grasp evaluation module, the task success rate is significantly improved compared to PointNetGPD in all settings. Specifically, the average task success rate is improved by $42.33\%$ in the 10-object setting, $13.16\%$ in the 20-object setting, and $30.33\%$ in the unseen-object setting. These results confirm that our sampling strategy can stably provide rich and effective grasp candidates, thereby improving the overall task success rate. However, the average grasp success rate decreases in all settings. We attribute this to the increased number of grasp attempts together with the limited accuracy of the PointNetGPD evaluation module. Notably, when the grasp candidate generation module of PointNetGPD is combined with our Stage-2 module, i.e., grasp evaluation and refinement, the average grasp success rate is substantially improved in all three settings. In particular, in the unseen-object setting, the grasp success rate is improved by $53.95\%$. Overall, the average task success rate is improved by $20.77\%$. These results strongly validate the accuracy, robustness, and generalization capability of our evaluation and refinement module. In addition, we observe that the refinement branch has only a limited effect on grasp success rate in the 10-object setting, but brings clear improvements in the more cluttered 20-object setting. Moreover, the average task success rate across the three settings is further improved by $6.94\%$, which verifies the necessity and effectiveness of the proposed refinement strategy.
@@ V-C Real-World Experiments
To further validate the generalizability of the proposed method, we construct a real-world grasping test set consisting of 30 similar objects and 25 novel objects whose shapes and sizes are substantially different from those in the training set. We design two cluttered real-world scenarios for evaluation: 1) randomly placing 10 similar objects and 2) randomly placing 10 novel objects. Each scenario is tested over 10 trials, and some representative scenarios are shown in Fig. 7. Specifically, we use an xMate ER7Pro robotic arm, an AG-105-145 gripper, and a ZED 2i camera. The entire system is integrated through ROS, and inverse kinematics is solved using MoveIt![4]. All methods employ FoundationStereo[28] to estimate depth from stereo RGB images and then use SAM[10] to segment the target object. For a fair comparison, all methods are evaluated under the same experimental setup and perception pipeline. The experimental results are reported in Table II.
@@
Chen et al.[3] 0.509 1.484 1.993
Ours 0.346 0.367 0.713
As shown in Table II, in the similar-object setting, our method achieves both a grasp success rate and a task success rate of $98\%$. Specifically, the average grasp success rate exceeds that of PointNetGPD by $39.18\%$ and that of Contact-GraspNet by $19.05\%$, while the average task success rate is higher than that of PointNetGPD by $78\%$ and that of Contact-GraspNet by $38\%$. In the novel-object setting, our method still maintains a solid performance, with both the grasp success rate and the task success rate remaining above $94\%$. In this setting, the average grasp success rate of our method is higher than that of PointNetGPD, Contact-GraspNet, and Chen et al. by $41.74\%$, $19.47\%$, and $24.89\%$, respectively, while the average task success rate is higher by $85.00\%$, $43.00\%$, and $9.00\%$, respectively. These results demonstrate the good generalization capability of the proposed method.
During the experiments, we also record the average time cost of each stage, as reported in Table III. Compared with the baselines, our method achieves the best real-time performance among all two-stage methods, with runtime only slightly higher than that of the end-to-end network Contact-GraspNet. In addition, we observe that the primary cause of grasp failure in real-world experiments is self-occlusion of the target object or occlusion caused by surrounding objects, which introduces unknown space in the scene and eventually causes collisions between the gripper and the object during execution. Furthermore, objects with severely non-convex geometries can also degrade the quality of grasp sampling, thereby increasing the likelihood of grasp failure.
@@ VI Conclusion
We proposed SuperGrasp, a new two-stage framework for single-view object grasping under incomplete geometric observations. It combines superquadric-based similarity matching for effective and diverse grasp candidate generation with an anchor-guided evaluation-refinement network for accurate and robust grasp evaluation and local refinement. Extensive experiments demonstrate the effectiveness of both stages and show that this combination makes single-view grasping more stable and efficient, while enabling solid performance and good generalization across diverse scenarios. Notably, all training data are collected entirely in simulation, and the model transfers directly to real-world scenes without any real-world fine-tuning.